# The Fragile Nature of Road Transportation Systems

### v2

@@ 5 Numerical simulation
Now we start the numerical simulation with different initial disruption demands $n^{\prime}$ from 1000 to 8000 vehicles. The simulation time is 7200 seconds for each scenario with different initial demands. Fig. 10(a) demonstrates that TTS grows exponentially with linearly increasing initial disruption demand, which validates the fragile nature proved with mathematical analysis. The solid, dashed, or dotted line each represents the TTS calculated under the three deterministic MFDs with green time being $\mu_{G}$, $\mu_{G}-\sigma_{G}$, or $\mu_{G}+\sigma_{G}$. Other than the black curves, there are also 1000 scattering points forming the blue curve. Each scatter point is composed of a full disruption recovery process sampled with equal intervals between an initial traffic demand from 1000 to 8000 vehicles. At each time step, a stochastic signal green time is chosen following the normal distribution $G\sim N(14.8,2.5)$ based on real-world data (Genser et al., 2024), leading to an uncertain MFD profile as Fig. 10(b) shows. This is an example of how the congestion is dissipated in the urban road network with an initial disruption of about 7000 vehicles, the green scattering points representing the trip completion at each time step are sampled from every $100^{th}$ point from the total 7200 time steps (seconds) for a clearer and more illustrative plotting.
@@
Since the blue curve composed from the scattering points closely aligns with the solid curve, it can be inferred that the influence of realistic stochasticity on the MFD is mostly negligible. Nevertheless, an intriguing observation is that, when the disruption demand is relatively low, the blue curve dips slightly below the MFD of the solid curve. However, the blue curve appears to exceed the TTS of the well-defined MFD when the demand is substantial. This may indicate that the recovery process with stochasticity can possibly have a larger second derivative (although two linear curves share the same second derivative of zero but different slopes can yield a similar observation). When we show the distribution of these two curves, as in Fig. 10(c), the TTS with stochasticity has a more concentrated distribution at a lower value while having a marginally longer tail pointing to the right, showing a more left-skewed distribution compared to the one without realistic stochasticity. This can also be validated by calculating the skewness of these two curves. When there is no stochasticity, the skewness is 0.67 while the skewness for the blue curve has a value of 0.70. As a greater skewness indicates a more fragile system, it means by introducing realistic stochasticity, the urban road network becomes even more fragile. It makes particular sense that as per definition, a fragile system should exhibit a much more degraded performance with larger disruptions brought by stochasticity, resulting in poor adaptability to uncertainties.
@@
Likewise, we showcase that supply disruptions can signify the fragile nature of transportation systems as well. With the same simulation environment, instead of the linearly increasing initial disruption demand, now a linearly growing supply disruption magnitude coefficient $r$ from 0 to 0.5 is considered. The simulation of the recovery process from the supply disruption along with 1000 uniformed sampled points between 0 and 0.5 is shown in Fig. 11(a). First of all, as the blue curve with stochasticity lies below the curve from the deterministic MFD with the mean green time, it means that the network with stochasticity has a better performance. This performance improvement can be attributed to the fact that, prior to reaching the maximal capacity, the upper MFD in the dotted curve generated from MoC keeps a larger space from the solid curve MFD profile compared to the distance between the lower MFD in the dashed curve. Therefore, although the likelihood of sampling a trip completion above or below the mean MFD profile is the same, there is a higher probability that the gained value of trip completion will surpass the loss caused by stochasticity. Despite this gain in system performance, when we calculate the skewness of distribution, we get a value of 0.49 for the deterministic MFD and 0.53 for the MFD when considering uncertainty, demonstrating again that stochasticity escalates the fragile response of the road transportation systems. Hence, the transportation systems have been demonstrated to be fragile with numerical simulation and such fragility has been reinforced with stochasticity in this work.
@@ 6 Conclusion
This research systematically demonstrated the fragile nature of road transportation systems with rigorous mathematical analysis and numerical simulation under realistic stochasticity. The mathematical proof comprehends the study of fragility under 1) microscopic - macroscopic, 2) demand disruption - supply disruption, and 3) onset of disruption - recovery from disruption. With essential assumptions regarding the disruption density in comparison to the critical density as well as the base traffic demand, we have validated the fragility of road transportation systems from various perspectives. Furthermore, through a numerical simulation with realistic data, we concluded that real-world stochasticity has a limited impact on the fragile characteristics of the system but contributes to rendering the system even more fragile. As early findings have also pointed out the possibilities of applying MFDs in other transportation modes, the fragility observed in urban road networks may well be extended to various transportation systems. This study aims to offer insights to researchers, emphasizing the fragile characteristics of transportation systems and encouraging the design of antifragile traffic control strategies in the future.

### v4

@@ 6.2 Demand disruptions
Now we start the numerical simulation with different initial disruption demands $n^{\prime}$ from 1000 to 8000 vehicles. The simulation time is 7200 seconds for each scenario with different initial demands. Fig. 13(a) demonstrates that TTS grows exponentially with linearly increasing initial disruption demand, which validates the fragile nature proved with mathematical analysis. The solid, dashed, or dotted line each represents the TTS calculated under the three deterministic MFDs with green time in $\mathcal{G}$. Other than the black curves, there are also 1000 scattering points forming the blue curve. Each scatter point is composed of a full disruption recovery process sampled following uniform distribution $n^{\prime}\sim U(1000,8000)$.
@@
Since the blue curve composed from the scattering points closely aligns with the solid curve, it can be inferred that the influence of realistic stochasticity on the MFD is mostly negligible. Nevertheless, an intriguing observation is that, when the disruption demand is relatively low, the blue curve dips slightly below the MFD of the solid curve. However, the blue curve appears to exceed the TTS of the well-defined MFD when the demand is substantial. This may indicate that the recovery process with stochasticity can possibly have a larger second derivative (although two linear curves share the same second derivative of zero but different slopes can yield a similar observation). When we show the distribution of these two curves, as in Fig. 13(b), the TTS with stochasticity has a more concentrated distribution at a lower value while having a marginally longer tail pointing to the right, showing a more left-skewed distribution compared to the one without realistic stochasticity. This can also be validated by calculating the skewness of these two curves. When there is no stochasticity, the skewness is 0.67 while the skewness for the blue curve has a value of 0.70. As a greater skewness indicates a more fragile system, it means by introducing realistic stochasticity, the urban road network becomes even more fragile. It makes particular sense that as per definition, a fragile system should exhibit a much more degraded performance with larger disruptions brought by stochasticity, resulting in poor adaptability to uncertainties.
6.3 Supply disruptions
@@
Likewise, we showcase that supply disruptions can strengthen the fragile nature of transportation systems as well. With the same simulation environment, instead of the linearly increasing initial disruption demand, now a linearly growing supply disruption magnitude coefficient $r$ from 0 to 0.5 is considered. The simulation of the recovery process from the supply disruption along with 1000 uniformed sampled points $r\sim U(0,0.5)$ is shown in Fig. 14(a). First of all, as the blue curve with stochasticity lies below the curve from the deterministic MFD with the mean green time, it means that the network with stochasticity has a better performance. This performance improvement can be attributed to the fact that, prior to reaching the maximal capacity, the upper MFD in the dotted curve generated from MoC keeps a larger space from the solid curve MFD profile compared to the distance between the lower MFD in the dashed curve. Therefore, although the likelihood of sampling a trip completion above or below the mean MFD profile is the same, there is a higher probability that the gained value of trip completion will surpass the loss caused by stochasticity. Despite this gain in system performance, when we calculate the skewness of distribution, we get a value of 0.49 for the deterministic MFD and 0.53 for the MFD with uncertainty, demonstrating again that stochasticity escalates the fragile response of the road transportation systems. Hence, the transportation systems have been demonstrated to be fragile with numerical simulation and such fragility has been reinforced with stochasticity in this work.
@@ 7 Conclusion
This research introduces the pioneering concept of (anti-)fragility and its detection under the context of transportation. Then it systematically demonstrates the fragile nature of road transportation systems through rigorous mathematical analysis. With established traffic models, namely FDs and MFDs, as the representation of the system dynamics, the second derivative of the system performance loss over the magnitudes of disruptions can be proven to be positive, indicating the fragile property. To achieve a universal conclusion, the fragility of the road transportation systems is validated from multiple perspectives: at the link level or the network level; under disruption from the demand or the supply side; and for the onset of disruption or the recovery from disruption. With such fragile nature established, this research also proposes a generic approach for quantifying and assessing network fragility by using a scalable unit MFD and a skewness-based indicator. An approximation function inspired by the Sigmoid curve has been developed for accurately determining the skewness, enabling the approximation and cross-comparison of the fragility of different networks using merely MFD-related parameters at hand. The proposed fragility indicator can be potentially applied to evaluate the fragility of future new infrastructures and transportation policies. Additionally, through a numerical simulation with realistic data, including topology attributes, driving behavior, signalization, etc., results suggest that real-world stochasticity has a limited effect on the fragile characteristics of road transportation systems but contributes to rendering the system even more fragile.
Several limitations of this study need to be acknowledged. Although uncertainties are explicitly accounted for in the numerical simulation, this research largely relies on the assumptions of homogenous networks and well-defined MFDs as in many other research works (Haddad and Mirkin, 2017; Zhou and Gayah, 2021; Genser and Kouvelas, 2022). Since hysteresis is a commonly observed phenomenon due to unevenly distributed traffic congestion, it can be of great interest to involve heterogeneity and hysteresis in future studies of antifragility. As hysteresis leads to a decrease in network serviceability, it can potentially be regarded as a virtual supply disruption. Additionally in Section 5, it is assumed that the backward wave is slower than the free flow speed, allowing for a relatively accurate skewness estimation with the proposed approximation function, but as noted, this assumption may be challenged by technological advancements, particularly with the rise of CAVs, highlighting the necessity for developing more precise approximation functions to address future challenges.
This study aims to offer insights to researchers, emphasizing the fragile nature of road transportation systems. Potential extensions of this work can be multifold. For instance, this study lays a theoretical foundation and complements our ongoing work on data-driven antifragile traffic control, countering the intrinsic fragility of road networks through induced antifragility. Furthermore, as early findings have pointed out the possibilities of applying MFDs in other transportation modes, the fragility demonstrated in urban road networks may well be extended to various transportation systems or even systems in other disciplines with similar characteristics. Given that network performance assessment in modern times should not rely on a sole criterion based on efficiency, a multi-objective framework should be developed in the future, incorporating factors such as efficiency, antifragility, sustainability, safety, and more.